# Dynamic Mixture of Progressive Parameter-Efficient Expert Library for Lifelong Robot Learning

### v1

@@ 4.2 Low-rank Expert Library
Following the pretrain-then-finetune paradigm, we pretrain the policy on a large-scale robotic dataset and freeze it during the lifelong adaptation to unseen downstream tasks.
Low-Rank Adaptation (LoRA) [36].
@@
A representative parameter-efficient fine-tuning (PEFT) method that introduces learnable low-rank matrices $\boldsymbol{A}\in\mathbb{R}^{d_{\text{in}}\times r}$, $\boldsymbol{B}\in\mathbb{R}^{r\times d_{\text{out}}}$ and integrate in parallel with the frozen weight matrix $\boldsymbol{W}_{0}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$ through addition, i.e., $\boldsymbol{W}_{0}+\boldsymbol{A}\boldsymbol{B}$. By leveraging the low-rank decomposition $r\ll\min\{d_{\text{in}},d_{\text{out}}\}$, LoRA substantially reduces the number of trainable parameters.
@@
Existing PEFT methods learn a task-specific low-rank adapter $\Delta\boldsymbol{W}_{k}=\boldsymbol{A}_{k}\boldsymbol{B}_{k}$ for each task $\mathcal{T}_{k}$ in lifelong robot learning [40, 41]. Such design requires task identifier during inference, lacks flexibility due to fixed capacity, and also hinders knowledge sharing across tasks. Prior arts in vision and language domain, e.g., LoRAHub [77] and SD-LORA [90], use a task-wise learnable weight vector to combine existing low-rank experts for better performance on the new task. A key insight is that these low-rank experts captures complementary knowledge from a variety of specific tasks and can be leveraged to solve a new task. However, lifelong learning in robotic domain could be even more challenging. First, compared to image or text classification [1, 3, 5] that requires only single forward inference, robotic manipulation is a sequential decision-making process with long horizon. Besides, robot learning requires a mixed types of knowledge, for example, visual concepts, textual task goals, different spatial relationships, and successful adaptation to a new task requires a combination of adaptation in each sub-modules.
Following this motivation, we propose to learn a progressive low-rank expert library on top of the pretrained policy and dynamically reuse low-rank experts from prior tasks $j\leq k-1$ to facilitate efficient forward transfer to new task $\mathcal{T}_{k}$. We equip each chosen pretrained linear layer =$\mathcal{F}=\{\boldsymbol{W}_{0},\boldsymbol{b}_{0}\}$ with a low-rank expert library =$\mathcal{L}=\{\boldsymbol{A}_{j},\boldsymbol{B}_{j},\boldsymbol{b}_{j}\}_{j=1}^{k}$ and use a global lightweight router $\mathcal{R}$ to dynamically integrate pretrained and task-specific knowledge based on the current context, providing a more flexible solution for adaptation.
@@
Specifically, the router $\mathcal{R}:\mathbb{R}^{d_{r}}\rightarrow\mathbb{R}^{6\times k}$ is implemented as a multi-layer perceptron (MLP) to obtain the dynamic coefficient for each low-rank expert. The router input $\boldsymbol{r}_{t}=[f^{v}_{t-T:t},f^{l},p_{t-T:t}]\in\mathbb{R}^{d_{r}}$, where $\cdot_{t-T:t}$ indicates mean pooling over the $T$-step context window. This aggregated representation contains information streams from the visual embeddings, the task instruction embeddings, and also the low-dimensional proprioceptive states, and hence comprehensively encodes the current context. Then we can dynamically obtain the expert coefficient:
$\boldsymbol{c}_{t}=[\boldsymbol{c}^{v}_{t};\boldsymbol{c}^{l}_{t};\boldsymbol{c}^{s}_{t};\boldsymbol{c}^{f}_{t};\boldsymbol{c}^{t}_{t};\boldsymbol{c}^{h}_{t}]=\text{top-k}(\mathcal{R}(\boldsymbol{r}_{t}),\delta),\quad\boldsymbol{c}_{t}^{\cdot}=[c_{1,t}^{\cdot},c_{2,t}^{\cdot},\cdots,c_{k,t}^{\cdot}]\in\mathbb{R}^{k}.$ (5)
[…]
$\boldsymbol{y}=\tilde{\boldsymbol{W}}\boldsymbol{x}+\tilde{\boldsymbol{b}}=\underbrace{\boldsymbol{W}_{0}\boldsymbol{x}+\boldsymbol{b}_{0}}_{\text{pretrained}}+\underbrace{\sum_{j=1}^{k}c_{j,t}\boldsymbol{A}_{j}\sum_{j=1}^{k}c_{j,t}\boldsymbol{B}_{j}\boldsymbol{x}+\sum_{j=1}^{k}c_{j,t}\boldsymbol{b}_{j}}_{\text{fine-tuned}}.$ (6)
@@
Upon the completion of task $\mathcal{T}_{k}$, we freeze all low-rank experts in the library as acquired knowledge and initialize new learnable expert for task $\mathcal{T}_{k+1}$. In order to reduce the interference between tasks, we employ the Gram-Schmidt orthogonalization method on $\boldsymbol{A}_{k+1}$ [59, 91], while other elements $\boldsymbol{B}_{k+1}$ and $\boldsymbol{b}_{k+1}$ are simply zero-initialized.
@@ 4.3 Expert Coefficient Replay
To mitigate catastrophic forgetting, we freeze the learned low-rank expert module after each task’s completion for future retrieval. However, since the router is continuously updated on sequentially arriving tasks, incorrect integration of the expert, i.e., the expert coefficients are inaccurate, may lead to severe policy distortion and suboptimal performance.
Inspired by replay methods [22, 24, 76], we propose an expert coefficient replay mechanism specifically designed for the lightweight router. The implementation involves two phases: During task finalization, we archive a subset of the router’s input-output pairs, i.e., the context embedding $\boldsymbol{r}$ and the coefficient vector $\boldsymbol{c}$, in a dedicated buffer $\mathcal{B}_{\text{CR}}$. When adapting to new tasks, the core idea is to enforce consistency by requiring the router to generate coefficients that closely match the stored historical coefficients given the same context. This is achieved through minimizing a mean squared error (MSE) objective:
$L_{\text{CR}}(\mathcal{R})=\mathbb{E}_{{(\boldsymbol{r}_{i},\boldsymbol{c}_{i,\text{old}})\sim\mathcal{B}_{\text{CR}}}}\frac{1}{2}\left(\mathcal{R}(\boldsymbol{r}_{i})-\boldsymbol{c}_{i,\text{old}}\right)^{2}.$ (7)
Compared with conventional expert demonstration replay that requires computation through the entire policy, our coefficient replay strategy demonstrates superior efficiency by focusing exclusively on the lightweight router and its inputs and outputs, yielding substantial reductions in both computational overhead and storage requirements.
@@ 5.1 Experimental Setup
We conduct evaluations in LIBERO [43], a lifelong learning benchmark on robot manipulation tasks. The robot in the benchmark is situated in a tabletop environment and equipped with a 6-DOF arm and a parallel gripper. We use four lifelong learning task suites: Goal, Spatial, Object, and Long. Each suite consists of 10 robotic manipulation tasks designed to investigate the transfer of knowledge related to task goals, spatial information, and various objects. Details can be found in Appendix A.3. The benchmark also includes LIBERO-90, a collection of 90 short-horizon diverse tasks that serves as a pretraining dataset for downstream transfer.
@@
We consider a naive sequential fine-tuning baseline, three representative lifelong learning methods that involve tuning the entire policy, and several state-of-the-art LoRA-based methods:
  • Sequential Fine-tuning (SeqFT) [43]: Naively perform full fine-tuning (FFT) or use frozen pretrained feature (FPF), i.e., freeze the spatial encoders and fine-tune the rest, on sequentially arriving robot learning tasks.
  • Experience Replay (ER) [22], a replay method that maintains a memory buffer with samples from previous tasks. We store a subset of data after each task and uniformly sample a fixed number of replay data from the buffer along with new task data during training.
  • Elastic Weight Consolidation (EWC) [27], a regularization method that uses the Fisher Information Matrix (FIM) to quantify the importance of each parameter and adds a regularization term to the learning objective to constrain network updates.
  • PackNet [30], an architectural method that first trains and prunes the network to retain a fixed proportion of the most important parameters. The selected parameters are then fine-tuned and frozen to prevent forgetting.
  • LifelOng knowledge Transfer Using Skills (LOTUS) [92] involves two stages: extracting temporal segment features by leveraging frozen foundation models for skill discovery, and hierarchical imitation learning with the meta-controller and a growing skill library.
  • Task-specific Adapters for Imitation Learning (TAIL) [40] introduces separate LoRAs into the spatial encoder and temporal decoder for each new task. The fusion module, state encoder, and action head are also full fine-tuned and stored as task-specific adapters. TAIL requires an oracle task identifier to retrieve the corresponding adapters during inference. In order to align with other methods, we also provide the result of SeqFT (LoRA), which fine-tunes LoRA parameters sequentially.
  • Learning to Modulate (L2M) [41] maintains a learnable modulation pool with keys and associated modulators. The embedded history of state tokens serves as the query vector for query-key matching, which finally steers the pretrained policy’s behavior. We follow the original paper to use a frozen fusion module and action head, and insert LoRA pools in the CLIP encoders and the temporal transformer backbone.
  • Incremental Learning of Retrievable Skills for Continual Imitation Learning (IsCiL) [42] establishes multifaceted prototypes through K-means clustering in the state space, where proper LoRA-based skills is retrieved upon current input state. We define the task-specific skill the same as in TAIL [40], i.e., LoRA matrices for transformer, the fusion module, the state encoder, and the action head.
@@
We utilize the open-source CLIP ViT/B-16 model [15] as vision and text encoders, while all other components of the policy is learned from robot demonstrations. We use a six-layer temporal transformer, using the same embedding size as the CLIP model $d=768$. The LoRA rank is set to 8 for CLIP encoders and 16 for others, the same as in [40]. We pretrain the base policy on the LIBERO-90 dataset and use it as the starting point of lifelong learning. During adaptation, we learn 10 epochs on each arriving task and evaluate every 2 epochs, using a batch size of 32 and AdamW optimizer with a learning rate of 1e-4. We perform expert coefficient replay for 10 epochs after each task. As for the expert router, we employ a scaled sigmoid output activation to restrict the coefficient in $[0,2]$ and use a top-3 strategy to select useful experts. All results are averaged over three random seeds. More details can be found in Appendix A.
Figure 3: Performance of the pretrained policy on LIBERO
@@ 5.2 Results and Analysis
Figure 5(a) examines the effects of various algorithm settings. Utilizing the top-3 experts yields better performance compared to using only the top-1 expert or merging all experts. Additionally, orthogonal initialization of new experts helps to reduce interference. Besides, directly synthesizing policy parameters and assigning distinct coefficients to different submodules allows for more flexible and effective integration of knowledge. Figure 5(b) illustrates the impact of the coefficient replay (CR) ratio on overall performance. CR ratio refers to the proportion of router input-output pairs stored in the buffer for future replay. The results show that coefficient replay does not hinder forward transfer but helps mitigate forgetting efficiently, even when only a small fraction (e.g., 5%) of previous router input-output pairs is retained. Additionally, we demonstrate that replaying the context embeddings and coefficient vectors incurs a much lower storage overhead (approx. 6.7%) compared to replaying full expert demonstrations.
Figure 6: Visualization of expert coefficients on the LIBERO-Object benchmark.
@@
DMPEL introduces two additional computational stages: (1) encoding the current context with the frozen backbone and using the global router to compute routing coefficients, and (2) averaging the parameters of the top-$\delta$ experts to synthesize $\tilde{\boldsymbol{W}}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$, which incurs approximately $2\times\delta\times r\times(d_{\text{in}}+d_{\text{out}})$ FLOPs. Thereafter, the forward pass through the linear layer requires the usual $2\times d_{\text{in}}\times d_{\text{out}}$ FLOPs. In robotic manipulation, which typically spans hundreds of decision steps and whose observations change only gradually from one step to the next, we conjecture that re-synthesizing policy parameters every several steps can yield a favorable trade-off between computational cost and task performance. Figure 5(c) verifies that, as the expert synthesize interval increases from 1 to 5, the average inference time rapidly approaches that of the backbone-only baseline (approx. 30 ms), while the success rate suffers only a slight decline. The result also indicates that the expert router indeed dynamically selects low-rank experts tailored to the current context.
Visualization on Expert Coefficients.
@@
We present a visualization of the expert coefficients for each sub-module during the lifelong learning process on the LIBERO-Object benchmark. As illustrated in Figure 6, significant knowledge sharing occurs across tasks, which contributes to the enhanced forward transfer capability of DMPEL. Notably, we observe that the low-rank expert learned for the action head in the first task is extensively reused in subsequent tasks, while the experts introduced later are rarely activated. This suggests that these inactive experts can be pruned to further improve storage efficiency without adversely affecting performance.
6 Conclusion
In this study, we introduce Dynamic Mixture of Progressive Parameter-Efficient Expert Library (DMPEL) for lifelong robot learning. Our approach builds an low-rank expert library incrementally and utilizes a lightweight router to dynamically integrate these experts. This enables the robot to adapt flexibly to various scenarios over its lifetime. Additionally, by utilizing the modular nature of the policy, we address the issue of catastrophic forgetting through coefficient replay on the router, which allows for precise retrieval of experts on previous tasks. This method is significantly more efficient than the expensive demonstration replay on the entire policy. Our extensive experiments on the LIBERO benchmark show that DMPEL surpasses state-of-the-art lifelong learning baselines in terms of forward transfer and catastrophic forgetting while requiring only very few trainable parameters and storage.
Limitations.
We conduct all experiments using a rather small model and we only use in-domain data to do pretraining and adaptation. We plan to further scale up our policy or employ more advanced VLA models (over billions of parameters) as the backbone in the future. In addition, we seek to further analyze the mechanisms of knowledge transfer and to enhance lifelong learning algorithms with a solid theoretical foundation.
Appendix A Experimental Details

### v2

@@ 4.2 Low-rank Expert Library
Following the pretrain-then-finetune paradigm, we pretrain the policy on a large-scale robotic dataset and freeze it during the lifelong adaptation to unseen downstream tasks. Low-Rank Adaptation (LoRA) (Hu et al., 2021) is a representative PEFT method that introduces learnable low-rank matrices $\bm{A}\in\mathbb{R}^{d_{\text{in}}\times r}$, $\bm{B}\in\mathbb{R}^{r\times d_{\text{out}}}$ and integrate in parallel with the frozen weight matrix $\bm{W}_{0}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$ through addition, i.e., $\bm{W}_{0}+\bm{A}\bm{B}$. By leveraging the low-rank decomposition $r\ll\min\{d_{\text{in}},d_{\text{out}}\}$, LoRA substantially reduces the number of trainable parameters.
@@
Existing methods usually learn a task-specific low-rank adapter $\Delta\bm{W}_{k}=\bm{A}_{k}\bm{B}_{k}$ for each task $\mathcal{T}_{k}$ in lifelong robot learning (Liu et al., 2024b; Schmied et al., 2023). Such design requires task identifier during inference, lacks flexibility due to fixed capacity, and also hinders knowledge sharing across tasks. Prior arts in vision and language domain, e.g., LoRAHub (Huang et al., 2024) and SD-LORA (Wu et al., 2025), use a task-wise learnable weight vector to combine existing low-rank experts for better performance on the new task. A key insight is that these low-rank experts captures complementary knowledge from a variety of specific tasks and can be leveraged to solve a new task. However, lifelong learning in robotic domain could be even more challenging. First, compared to image or text classification (De Lange et al., 2021; Masana et al., 2022; Wang et al., 2024a) that requires only single forward inference, robotic manipulation is a sequential decision-making process with long horizon. Besides, robot learning requires a mixed types of knowledge, for example, visual concepts, textual task goals, different spatial relationships, and successful adaptation to a new task requires a combination of adaptation in each sub-modules.
@@
Following this motivation, we propose to learn a progressive low-rank expert library on top of the pretrained policy and dynamically reuse low-rank experts from prior tasks to facilitate efficient forward transfer to new task $\mathcal{T}_{k}$. We equip each chosen pretrained linear layer $\mathcal{F}=\{\bm{W}_{0},\bm{b}_{0}\}$ with a low-rank expert library $\mathcal{L}=\{\bm{A}_{j},\bm{B}_{j},\bm{b}_{j}\}_{j=1}^{k}$ and use a lightweight router $\mathcal{R}$ to dynamically integrate pretrained and task-specific knowledge based on the current context, providing a more flexible solution for adaptation. Specifically, the router $\mathcal{R}:\mathbb{R}^{d_{r}}\rightarrow\mathbb{R}^{M\times k}$ is implemented as a multi-layer perceptron (MLP) to obtain the dynamic coefficient for each low-rank expert. The router input $\bm{r}_{t-T:t}=[\tilde{f}^{v}_{t-T:t},\tilde{f}^{l}_{t-T:t},p_{t-T:t}]\in\mathbb{R}^{d_{r}}$, where $\cdot_{t-T:t}$ indicates mean pooling over the $T$-step context window. This aggregated representation contains information streams from the visual embeddings, the task instruction embeddings, and also the low-dimensional proprioceptive states, and hence comprehensively encodes the current context. Then we can obtain the expert coefficient:
$\bm{c}_{t}=\text{top-k}(\mathcal{R}(\bm{r}_{t-T:t}),\delta)\in\mathbb{R}^{M\times k}.$ (5)
Note that the output $\bm{c_{t}}$ is divided into $M$ distinct expert coefficient vectors $\bm{c}_{t}^{\diamond}=[c_{1,t}^{\diamond},c_{2,t}^{\diamond},\cdots,c_{k,t}^{\diamond}]\in\mathbb{R}^{k}$, each assigned to a specific sub-module $\diamond$ in the policy. In our case, we have $M=6$, consisting of vision encoder, text encoder, state encoder, modality fusion module, temporal transformer, and action head. This design allows each sub-module to use a different coefficient vector to integrate knowledge, providing larger flexibility in adaptation. The $\text{top-k}(\cdot,\delta)$ operator retains only the top $\delta$ over $k$ entries of the input vector at their original values, while setting all other coefficients to zero. This ensures efficient utilization of experts by reducing computational overhead while maintaining flexibility for adaptation. Finally, the input-conditioned linear layer $\{\tilde{\bm{W}},\tilde{\bm{b}}\}$ is the dynamic mixture of the pretrained weight $\mathcal{F}$ and the low-rank experts from the library $\mathcal{L}$:
$\bm{y}$ $=\bm{x}\tilde{\bm{W}}+\tilde{\bm{b}}=\underbrace{\bm{x}\bm{W}_{0}+\bm{b}_{0}}_{\text{pretrained}}+\underbrace{\bm{x}\sum_{j=1}^{k}c_{j,t}\bm{A}_{j}\sum_{j=1}^{k}c_{j,t}\bm{B}_{j}+\sum_{j=1}^{k}c_{j,t}\bm{b}_{j}}_{\text{fine-tuned}}.$ (6)
@@
Upon the completion of task $\mathcal{T}_{k}$, we freeze all low-rank experts in the library as acquired knowledge and initialize new learnable expert for task $\mathcal{T}_{k+1}$. In order to reduce the interference between tasks, we employ the Gram-Schmidt orthogonalization method on $\bm{A}_{k+1}$ (Smith et al., 2023; Wang et al., 2023a), while other elements $\bm{B}_{k+1}$ and $\bm{b}_{k+1}$ are simply zero-initialized.
@@ 4.3 Expert Coefficient Replay
To mitigate catastrophic forgetting, we freeze the learned low-rank expert after each task’s completion for future retrieval. However, since the router is continuously updated on sequentially arriving tasks, incorrect integration of the expert, i.e., the expert coefficients are inaccurate, may lead to severe policy distortion and suboptimal performance.
Inspired by replay methods (Chaudhry et al., 2019; Xie & Finn, 2022; Zhao et al., 2024), we propose an expert coefficient replay mechanism to regularize the lightweight router on previously encountered tasks. The implementation involves two phases: During task finalization, we archive a subset of the router’s input-output pairs, i.e., the context embedding $\bm{r}$ and the coefficient vector $\bm{c}$, in a dedicated buffer $\mathcal{B}_{\text{CR}}$. When adapting to new tasks, the core idea is to enforce consistency by requiring the router to generate coefficients that closely match the stored historical coefficients given the same context, which can be achieved through minimizing a mean squared error (MSE) objective:
$L_{\text{CR}}(\mathcal{R})=\mathbb{E}_{{(\bm{r}_{i},\bm{c}_{i,\text{old}})\sim\mathcal{B}_{\text{CR}}}}\frac{1}{2}\left(\mathcal{R}(\bm{r}_{i})-\bm{c}_{i,\text{old}}\right)^{2}.$ (7)
Compared to conventional experience replay methods that require storing entire trajectories and processing them through the entire policy, our expert coefficient replay strategy leverages the modularity of the low-rank expert library and focuses on regularizing the lightweight router using low-dimensional embeddings and coefficients, resulting in significant reductions in both computational overhead and storage requirements.
@@ 5.1 Experimental Setup
We present an overview of the experimental setup below, with details included in Appendix A.
@@
We conduct extensive evaluation in a lifelong robot learning benchmark LIBERO (Liu et al., 2024a). The robot is situated in a tabletop environment and equipped with a 6-DOF arm and a parallel gripper. We use four lifelong learning task suites: Goal, Spatial, Object, and Long. Each suite consists of 10 sequentially arriving robotic manipulation tasks designed to investigate the transfer of knowledge related to task goals, spatial information, and various objects. The benchmark also includes LIBERO-90, a collection of 90 short-horizon diverse tasks that serves as a pretraining dataset for downstream transfer. All tasks are consistent with the formulation in Section 3, i.e., generating continuous actions to control the robot based on multi-modal input.
@@
We consider sequential fine-tuning (SeqFT) baselines (Liu et al., 2024a) that either perform full fine-tuning (FFT) or use frozen pretrained feature (FPF) naively on sequentially arriving tasks. We also consider three representative lifelong learning methods that fine-tune all parameters, including Experience Replay (ER) (Chaudhry et al., 2019), Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017), and PackNet (Mallya & Lazebnik, 2018). We also include a hierarchical baseline, LifelOng knowledge Transfer Using Skills (LOTUS) (Wan et al., 2024). Besides, we consider several LoRA-based methods that use different low-rank expert inserting and retrieving mechanism, including Task-specific Adapters for Imitation Learning (TAIL) (Liu et al., 2024b), Learning to Modulate (L2M) (Schmied et al., 2023), and Incremental Learning of Retrievable Skills for Continual Imitation Learning (IsCiL) (Lee et al., 2024). Since TAIL requires oracle task identifiers, we also provide the result of SeqFT (LoRA) to align with other methods.
@@
We utilize CLIP (Radford et al., 2021) as vision and text encoders, while all other components of the policy is learned from robot demonstrations. The LoRA rank is set to 8 for CLIP encoders and 16 for others, the same as in (Liu et al., 2024b). We pretrain the base policy on the LIBERO-90 dataset and use it as the starting point of lifelong learning. During adaptation, we learn 10 epochs on each arriving task and evaluate every 2 epochs, using a batch size of 32 and AdamW optimizer with a learning rate of 1e-4. We perform expert coefficient replay for 10 epochs after each task. As for the expert router, we employ a scaled sigmoid output activation to restrict the coefficient in [0,2] and use a top-3 strategy to select useful experts. All results are averaged over three random seeds.
@@ 5.2 Results and Analysis
Figure 5(a) examines the effects of various algorithm settings. Utilizing the top-3 experts yields better performance compared to using only the top-1 expert or merging all experts. Besides, assigning distinct coefficients to different submodules allows for more flexible and effective integration of knowledge. Figure 5(b) illustrates the impact of the coefficient replay (CR) ratio on overall performance. CR ratio refers to the proportion of router input-output pairs stored in the buffer for future replay. The results show that coefficient replay does not hinder forward transfer but helps mitigate forgetting efficiently, even when only a small fraction (e.g., 5%) of previous router input-output pairs is retained. Additionally, we demonstrate that replaying the context embeddings and coefficient vectors incurs a much lower storage overhead (approx. 6.7%) compared to replaying full expert demonstrations.
((a))
((b))
Figure 6: Analysis on expert coefficients. (a) Visualization of coefficients for each sub-module in the policy on LIBERO-Object (b) Influence on success rate (SR) when pruning underactivated experts.
@@
DMPEL introduces two additional stages: (1) encoding the current context with the frozen backbone and using the global router to compute routing coefficients, and (2) averaging the parameters of the top-$\delta$ experts to synthesize $\tilde{\bm{W}}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$, which incurs approximately $2\times\delta\times r\times(d_{\text{in}}+d_{\text{out}})$ FLOPs. Thereafter, the forward pass through the linear layer requires the usual $2\times d_{\text{in}}\times d_{\text{out}}$ FLOPs. In robotic manipulation, which typically spans hundreds of decision steps and whose observations change only gradually from one step to the next, we conjecture that re-synthesizing policy parameters every several steps can yield a favorable trade-off between computational cost and task performance. Figure 5(c) verifies that, as the expert synthesize interval increases from 1 to 5, the average inference time rapidly approaches that of the backbone-only baseline (approx. 30 ms), while the success rate suffers only a slight decline. The result also indicates that the router indeed dynamically selects low-rank experts tailored to the current context.
Figure 7: t-SNE visualization of embeddings from the final block when evaluated on Task 2 from LIBERO-Object.
Visualization Analysis.
@@
We present a visualization of the expert coefficients for each sub-module during the lifelong learning process on the LIBERO-Object benchmark. As illustrated in Figure 6(a), significant knowledge sharing occurs across tasks, which contributes to the enhanced forward transfer capability of DMPEL. Notably, we observe that the low-rank expert learned for the action head in the first task is extensively reused in subsequent tasks, while the experts introduced later are rarely activated. This suggests that these inactive experts can be pruned to further improve storage efficiency without adversely affecting performance, as shown in Figure 6(b). We also present t-SNE visualization in Figure 7, illustrating the latent embeddings from the temporal transformer’s final block during the evaluation on Task 2. DMPEL maintains embedding consistency after adapting to new tasks, whereas SeqFT results in substantial representation drift and catastrophic forgetting.
Due to page limit, we provide additional results in Appendix B, including further ablation studies, visualization of trajectories, and lifelong learning curves.
6 Conclusion and Limitation
We present Dynamic Mixture of Progressive Parameter-Efficient Expert Library (DMPEL), a lifelong robot learning framework that incrementally builds a low-rank expert library and uses a lightweight router with coefficient replay to dynamically combine experts, achieving high forward transfer and low catastrophic forgetting with high efficiency in both storage and computation. Experiments on the LIBERO benchmark demonstrate that DMPEL achieves superior performance compared to state-of-the-art baselines. However, currently all experiments employ a rather small model with in-domain data for pretraining and adaptation. Future work will explore scaling up to larger and advanced policies (e.g., VLA models with over billions of parameters) and transferring to real robots, as well as deepen the theoretical understanding of knowledge transfer and retention in lifelong learning.
Reproducibility Statement
Our work is reproducible using the code available at https://github.com/HarryLui98/DMPEL. Additionally, we include detailed pseudocode and hyperparameter settings in the Appendix.
Appendix A Experimental Setup